OpenStudy (anonymous):

explain the expression bellow and how to get the form : sin a + sin b = 2 [sin (a+b)/2] x [cos (a-b)/2]

OpenStudy (apoorvk):

Do you need the derivation of the formula?

OpenStudy (anonymous):

yes i need ,,,, help please

OpenStudy (apoorvk):

Let A = (U+V) and let B = (U-V) so, we have: sin(U+V) + sin(U-V) =sinUcosV + cosUsinV + sinUcosV - cosUsinV =2sinUcosV Now, form above, A+B = U + V + U - V =2U, and similarly, A-B = 2V so, U = (A+B)/2 and V=(A-B)/2 hence, the equation = 2sin((A+B)/2)cos((A-B)/2).